OpenStudy (anonymous):

graph the linear equation y= 2/3 (x+9)

OpenStudy (anonymous):

@mathstudent55 ??

OpenStudy (anonymous):

@AccidentalAiChan ??

OpenStudy (accidentalaichan):

The only advice I may give you is: H- Horizontal O- Means zero slope Y- equal to Y V- Vertical U- Undefined slope X- equal to x Hope this helps :3

OpenStudy (accidentalaichan):

I think the coordinates are (3, -3) and (-3, -3)

OpenStudy (anonymous):

Have you tried expanding the bracket (right side)?

OpenStudy (anonymous):

if you do that then you will get the following \[y=\frac{ 2 }{ 3 }x +6\]This equation is now in the standard form of a linear graph y=mx+c, where 'y' & 'x' are input and output(i.e. you input a value of 'x' and you get and output value which is 'y'), 'm' is the gradient (steepness) of the line and 'c' in the y-intercept.

OpenStudy (anonymous):

okay now graph it please @Mousam

OpenStudy (anonymous):

ok but first of all you need to know another point with is the x-intercept...

OpenStudy (anonymous):

to get that, you subsitute 0 as a value of y. Doing that will give you the following \[0=\frac{ 2 }{ 3 }x+6\]\[0=2x+18\]\[2x=-18\]\[x=-9\] which means that the line will intercept the x-axis at (-9,0)

OpenStudy (anonymous):

now that we know the x-intercept (-9,0) and y-intercept (0,6), we can graph the line easily. Just plot the two known intercept and join them together..

OpenStudy (anonymous):